OpenStudy (anonymous):

Devaughn's age is two times Sydney's age. The sum of their ages is 57 . What is Sydney's age?

OpenStudy (anonymous):

I got 14.5. Is this multi choice or an open ended question?

OpenStudy (anonymous):

57 / 2 = 28.5 / 2 = 14.5

OpenStudy (anonymous):

open ended

OpenStudy (anonymous):

Okay so i believe its 14.5. Unless you can't have a decimal answer. But you could round it to 15 if you wanted.

OpenStudy (anonymous):

it was 19